To open a clothing repair shop, you need to carry out the following expenses: purchase of raw materials and materials - 50 thousand UAH. per year, the cost of equipment - 200 thousand UAH. (service life - 10 years), rental of premises - 3 thousand UAH. monthly, salary of employees - UAH 8.4 thousand. monthly, electricity - 18 thousand UAH. in year.

For the amount of funds invested in production, it would be possible to receive an income of 7% per annum, as well as receive wages in a similar production - 30 thousand UAH. in year.

Calculate accounting, economic and normal profit if the total income of the enterprise is 550 thousand UAH. in year


1
Expert's answer
2021-03-04T14:55:21-0500

Answer:

Step 1


According to the above question for calculating the accounting profit we need to take the total income of the enterprise, purchase of raw materials and materials, the cost of equipment, rental of premises in a year, salary of employees in a year and the monthly electricity given in a year.

Similarly, for calculating the Economic Profit we need to take the desired result of the accounting profit, the purchase of raw materials and materials, the cost of equipment, rental of premises in a year, salary of employees in a year and the monthly electricity given in a year, rate percentage of income that is 7% and the wages received.

Here the normal profit is taken as zero.

 

 

Step 2

Therefore, the Accounting Profit is:

Given, the total income of the enterprise= 550 thousand UAH,

purchase of raw materials and materials = 50 thousand UAH, the cost of equipment= 200 thousand UAH, rental of premises = 3 thousand UAH, so for a year it would be (3"\times"12)thousand UAH = 36 thousand UAH, salary of employees = 8.4 thousand UAH, so for a year it would be (8.4"\times"12)thousand UAH= 100.8 thousand UAH, electricity= 18 thousand UAH.

So, the Accounting Profit= total income of the enterprise - (purchase of raw materials + the cost of equipment + rental premises in a year + salary of employees in a year + electricity)

 Accounting profit = [550-(50+200+36+100.8+18) ]thousand UAH

               = [550-404.8] thousand UAH

               =145.2 thousand UAH

Similarly, the Economic Profit is;

Given, the desired result of the accounting profit = 145.2 thousand UAH, the purchase of raw materials and materials = 50 thousand UAH, the cost of equipment = 200 thousand UAH, rental of premises in a year = 36 thousand UAH, salary of employees in a year = 100.8 thousand UAH ,the monthly electricity given in a year = 18 thousand UAH, rate percentage of income that is 7%, wages received = 30 thousand UAH

Economic Profit =[145.2-{(50+200+36+100.8+18)"\times""\times""7 \over 100"+30}]thousand UAH

             =[145.2-{404.8"\times""7 \over 100" +30}] thousand UAH

             = [145.2-58.336] thousand UAH

             = 86.864 thousand UAH

Normal Profit is: NP=0